EMISSION SPECTRA OF NiH SHOWING EXTENSIVE ELECTRONIC ENERGY TRANSFER BETWEEN EXCITED STATES

\underline{AMANDA J. ROSS}, RAPHAEL VALLON, PATRICK CROZET and CYRIL RICHARD, Université Lyon 1; CNRS; LASIM UMR 5579, 43 Bd du 11 novembre 1918, F-69622 Villeurbanne, France.

Fourier transform resolved laser induced fluorescence spectra have allowed us to investigate the low-lying electronic states of NiH arising from the Ni(3d^{9} 4s^{1}) configuration. Although laser excitation with a single mode laser should promote the molecule (formed in a hollow cathode discharge running at 1 torr pressure, with 10% H_{2} in argon) to a unique excited state that should be readily identifiable from the work of Kadavathu and co-workers,^{a} the resolved fluorescence spectra are unexpectedly complex. Efficient energy transfer occurs between excited electronic states, retaining only isotopic selectivity. Analysis of the spectra locates energy levels up to 8000 cm^{−1} above v=0 in the electronic ground state (X ^{2}Δ_{5/2}), from all components of the ^{2}Σ, ^{2}Π and ^{2}Δ states forming the supermultiplet complex studied by Gray et al.^{b} Measurements have been made so far for ^{58}NiH and ^{60}NiH. To observe heavier isotopes of nickel, an intracavity fluorescence experiment is required. Signals are increased by a factor 25, but the signal/noise ratio in Fourier spectra is disappointing when compared with intracavity excitation results.^{c}^{d} Nevertheless, this work has allowed us almost to double the range of known energy levels below 10000 cm^{−1} for NiH.

aS. A. Kadavathu, R. Scullman, R. W. Field, J. A. Gray, and M. Li, J. Mol. Spectrosc., \underline{147} 448 (1991)

bJ. A. Gray, M. Li, T. Nelis, and R. W. Field, J. Chem. Phys., 1991, \underline{95} 7164 (1991)

cE. J. Hill and R. W. Field, J. Chem. Phys., \underline{93} 1 (1990)

dL. C. O’Brien and J. J. O’Brien, Astrophys. J., \underline{621} 554 (2005)